# Did I Faithfully Say What I Thought? Bridging the Gap Between LLMs Neural Activity and Self-Explanations

### v3

@@ F NeuroFaith Faithfulness Measure Comparison
In this section we examine how faithfulness as measured by NeuroFaith compares to existing Counterfactual Intervention (CI) Atanasova et al. (2023) and Attribution Agreement (AA) approaches Parcalabescu & Frank (2024) for evaluating LLM self-NLE faithfulness. We focus on 2-hop reasoning and begin with a qualitative comparison that illustrates potential differences between these approaches, followed by a quantitative analysis that provides evidence for these observations.
@@
To illustrate the critical differences between existing faithfulness evaluation approaches and NeuroFaith, we examine a concrete 2-hop reasoning example (see Figure 32) that illusrrates fundamental limitations in current methods (CI and AA). Given input text $x$, prediction $f(x)$, and self-generated explanation $e(x)$, we compare how different faithfulness approaches evaluate the same case. The Example Setup:
  • Original input: $x$ = ”The father of Carol Chomsky is
  • Original prediction: $f(x)$ = ”Harry Abraham Schatz
@@
Counterfactual Intervention (CI) would assess the self-NLE $e(x_{cf})$ as faithfulness. This evaluation is based solely on the presence of the intervention term ”spouse” within $e(x_{cf})$, establishing consistency between the input modification and explanation content. Moreover, Attribution Analysis (AA) methods consists in comparing attribution scores (e.g., using SHAP) to highlight important tokens (e.g. ”father”, ”spouse”, ”Carol”, ”Chomsky”) for the prediction and the self-NLE. High correlation coefficients between the attribution vectors for prediction and self-NLE would similarly classify the self-NLE as faithful.
However, NeuroFaith rejects $e(x_{cf})$ as unfaithful because the bridge object (”Morris Halle”) is not contained in the decoded hidden states ($\{\tilde{h^{\ell}_{k}}\}$). This mismatch between internal computation and self-NLE content reveals unfaithfulness. The steered self-NLE obtained with linear faithful steering (see Section 6) is evaluated as faithful by NeuroFaith, due to shared bridge object (”Noam Chomsky”) between the new self-NLE and the decoded hidden states.
@@
We propose an experimental protocol to evaluate the practical utility of different faithfulness measures for model analysis. The protocol is motivated by two common applications of explanations in AI systems: (1) troubleshooting models by identifying the source of errors in wrong predictions Biecek & Samek (2024), and (2) detecting potential biases or shortcuts in correct predictions to ensure they align with expected reasoning processes Ribeiro et al. (2016).
@@
Each category can be further subdivided with respect to faithfulness, enabling comparison between faithful and unfaithful self-NLE within each reasoning pattern. In the following we propose 3 metrics based on these 4 categories and assess if faithful self-NLE lead to better model debugging and bias targeting than unfaithful self-NLE.
@@
$CAS(A,B,hint1)=log\Bigl(\frac{PR(A,B,hint1,faithful)}{PR(A,B,hint1,unfaithful)}\Bigr)$ (5)
A positive CAS indicates that faithful self-NLE better identify first-hop failures.
@@
NeuroFaith CI NeuroFaith CI NeuroFaith CI
hint1 0.04 -0.07 0.06 -1.10 0.75 -0.99
hint2 -0.44 -0.55 0.28 -0.03 -0.35 1.25
$r_{2}\rightarrow r^{\prime}_{2}$ 0.09 -2.62 0.17 -0.32 0.31 -0.26
Table 12: NeuroFaith comparison to CI (higher is better) across models on 352 CI-compatible samples on 2-hop reasoning.
@@
We now test whether faithful self-NLE better identify second-hop reasoning failures through a targeted intervention. Following the logic introduced above, for each input $x$ having led to a wrong prediction ($\widehat{o}_{3}\neq o_{3}$), we create a modified version $x_{hint2}=(o_{2},r_{2},o_{3},x)$ that explicitly provides the second part of the 2-hop reasoning. For example:
@@
We test whether faithful self-NLE better identify non-canonical reasoning pathways through the modification of the second step of the reasoning trace ($r_{2}$). Intuitively, a non-canonical reasoning pathways is more prone to lead to false reasoning when chaning one step of the 2-hop reasoning. For each input $x$ having led to a good prediction ($\widehat{o}_{3}=o_{3}$), we create a modified version $x_{r_{2}\rightarrow r^{\prime}_{2}}=(o_{1},r_{1},{\blacktriangle{}},r_{2}^{\prime},\bullet)$ that change the second relation of the 2-hop reasoning input. For example:
@@
If faithful self-NLE accurately reflect internal reasoning, then changing $r_{2}$ should differentially decrease performance for Category C (alternative reasoning pathway) versus Category D (canonical reasoning), and this difference should be stronger for faithful self-NLE. We build our third metric following the notations introduced above, based on the Compound Accuracy Score:
@@
We evaluate our three faithfulness indicators using both NeuroFaith and Counterfactual Intervention (CI). The Wikidata-2-hop dataset provides natural support for computing $x_{r_{2}\rightarrow r^{\prime}_{2}}$ instances, as it contains multiple reasoning chains involving similar objects. Additionally, the dataset includes counterfactual interventions through variations in reasoning chains, enabling straightforward CI computation across multiple instances.
Our evaluation reveals three key findings. First, Table 14 shows that NeuroFaith consistently outperforms CI on the subset of CI-compatible samples. Second, Table 14 presents NeuroFaith results on the complete dataset, showing positive metric values across all models except gemma-2-2b.
@@ G Classification Examples
In this section we give two examples of instances characterized as either faithful or unfaithful in classification case.
@@
Figure 36: Example from the Wikidata-2-hop dataset where we observe an internal-external reasoning mismach. The model incorrectly answers ”Jobete Music” and provides a self-NLE referencing to the bridge object ”Sam Cooke”. This bridge object is false and does not appear in the set of natural language interpretations of $f$ latent states. However, the ground truth bridge object ”Louis Armstrong” is decoded at token index 11 and layers 13 and 14.

### v4

@@ F NeuroFaith Faithfulness Measure Comparison
In this section we examine how faithfulness as measured by NeuroFaith compares to existing Counterfactual Intervention (CI) (Atanasova et al., 2023) and Attribution Agreement (AA) approaches (Parcalabescu and Frank, 2024) for evaluating LLM self-NLE faithfulness. We focus on 2-hop reasoning and begin with a qualitative comparison that illustrates potential differences between these approaches, followed by a quantitative analysis that provides evidence for these observations.
@@
To illustrate the critical differences between existing faithfulness evaluation approaches and NeuroFaith, we examine a concrete 2-hop reasoning example (see Figure 35) that illustrates fundamental limitations in current methods (CI and AA). Given input text $x$, prediction $f(x)$, and self-generated explanation $e(x)$, we compare how different faithfulness approaches evaluate the same case. We analyse the following example:
  • Input: $x$ = ”The father of Carol Chomsky is
  • Prediction: $f(x)$ = ”Harry Abraham Schatz
@@
Counterfactual Intervention (CI) would assess the self-NLE $e(x_{cf})$ as faithfulness. This evaluation is based solely on the presence of the intervention term ”spouse” within $e(x_{cf})$, establishing consistency between the input modification and explanation content. Attribution Analysis (AA) methods consists in comparing attribution scores (e.g., using SHAP) to highlight important tokens (e.g. ”father”, ”spouse”, ”Carol”, ”Chomsky”) for the prediction and the self-NLE. High correlation coefficients between the attribution vectors for prediction and self-NLE would similarly classify the self-NLE as faithful.
On the contrary, NeuroFaith rejects $e(x_{cf})$ as unfaithful because the bridge object (”Morris Halle”) is not contained in the decoded hidden states ($\{\tilde{h^{\ell}_{k}}\}$). This mismatch between internal computation and self-NLE content reveals unfaithfulness. The steered self-NLE obtained with linear faithful steering (see Section 6) is evaluated as faithful by NeuroFaith, due to shared bridge object (”Noam Chomsky”) between the new self-NLE and the decoded hidden states.
@@
We propose an experimental protocol to evaluate the practical utility of different faithfulness measures for model analysis. The protocol is motivated by two common applications of explanations in AI systems: (1) troubleshooting models by identifying the source of errors in wrong predictions (Biecek and Samek, 2024), and (2) detecting potential biases or shortcuts in correct predictions to ensure they align with expected reasoning processes (Ribeiro et al., 2016).
@@
Each category can be further subdivided with respect to faithfulness, enabling comparison between faithful and unfaithful self-NLE within each reasoning pattern. In the following we propose 3 evaluation protocols based on these 4 categories and assess if faithful self-NLE lead to better model debugging and bias targeting than unfaithful self-NLE.
@@
$CAS(A,B,hint1)=\log\Bigl(\frac{PR(A,B,hint1,faithful)}{PR(A,B,hint1,unfaithful)}\Bigr)$ (5)
A positive CAS indicates that faithful self-NLE better identify first-hop failures. This metric is conceptually close to the In-Context Editing instantiation of the approach proposed by Zaman and Srivastava (2025) to compare faithfulness measures. In the following, we extend beyond this proposal with two other metrics.
@@
NeuroFaith CI AA NeuroFaith CI AA NeuroFaith CI
hint1 0.04 -0.07 -0.23 0.06 -1.10 0.40 0.75 -0.99
hint2 -0.44 -0.55 -1.23 0.28 -0.03 0.15 -0.35 1.25
$r_{2}\rightarrow r^{\prime}_{2}$ 0.09 -2.62 1.32 0.17 -0.32 0.01 0.31 -0.26
Table 15: NeuroFaith comparison to CI and AA (higher is better) across models on 352 CI-compatible samples on 2-hop reasoning.
@@
We then test whether faithful self-NLE better identify second-hop reasoning failures through a targeted intervention. Following the logic introduced above, for each input $x$ having led to a wrong prediction ($\widehat{o}_{3}\neq o_{3}$), we create a modified version $x_{hint2}=(o_{2},r_{2},o_{3},x)$ that explicitly provides the second part of the 2-hop reasoning. For example:
@@
We test whether faithful self-NLE better identify non-canonical reasoning pathways through the modification of the second step of the reasoning trace ($r_{2}$). Intuitively, a non-canonical reasoning pathway is more prone to leading to false reasoning when changing one step of the 2-hop reasoning. For each input $x$ having led to a correct prediction ($\widehat{o}_{3}=o_{3}$), we create a modified version $x_{r_{2}\rightarrow r^{\prime}_{2}}=(o_{1},r_{1},\tiny{\blacktriangle{}},r_{2}^{\prime},\bullet)$ that changes the second relation of the 2-hop reasoning input. For example:
@@
If faithful self-NLE accurately reflect internal reasoning, then changing $r_{2}$ should significantly decrease performance for Category C (alternative reasoning pathway) versus Category D (canonical reasoning), and this difference should be stronger for faithful self-NLE. We build our third metric following the notations introduced above, based on the Compound Accuracy Score:
@@
We evaluate our three faithfulness indicators to compare NeuroFaith to commonly used Attribution Agreement (AA) and Counterfactual Intervention (CI) methods. The Wikidata-2-hop dataset provides natural support for computing $x_{r_{2}\rightarrow r^{\prime}_{2}}$ instances, as it contains multiple reasoning chains involving similar objects. Additionally, the dataset includes counterfactual interventions through variations in reasoning chains, enabling straightforward CI computation as in Atanasova et al. (2023) across multiple instances. We implement AA based on gradient-based attributions as in Wiegreffe et al. (2021). Due to prohibitive cost for 27-billion models, we only apply AA to Gemma-2-2B and Gemma-2-9B.
Our evaluation reveals 2 key findings. First, Table 16 shows NeuroFaith outperforms CI and AA on average on the subset of CI-compatible samples. Second, Table 16 presents NeuroFaith results on the complete dataset, showing positive metric values across all models except gemma-2-2b.
@@ G Classification Examples
This section gives two examples of instances characterized as either faithful or unfaithful in classification case.
@@
Figure 39: Example from the Wikidata-2-hop dataset where we observe an internal-external reasoning mismach. The model incorrectly answers ”Jobete Music” and provides a self-NLE referencing to the bridge object ”Sam Cooke”. This bridge object is false and does not appear in the set of natural language interpretations of $f$ latent states. However, the ground truth bridge object ”Louis Armstrong” is decoded at token index $1$ and layers 13 and 14.